OpenStudy (stickman1998):

The volume in cubic feet of a box can be expressed as, (x)=x^3 - 6x^2+8x, or as the product of three linear factors with integer coefficients. The width of the box is x-2. Factor the polynomial to find linear expressions for the height and length. Show your work

OpenStudy (math&ing001):

You can start by factoring x f(x)=x^3 - 6x^2 + 8x = x(x^2-6x+8) From here you know that one other factor is the width (x-2), so we'll use that. f(x) = x(x^2-2x-4x+8) = x(x(x-2)-4(x-2)) = x(x-2)(x-4) I'll let you figure out which is the height and the length.

OpenStudy (stickman1998):

Wow that's a lot of parenthesis, I'm not sure which would be the height and length?

OpenStudy (math&ing001):

Well remember that the length is bigger that the width ;]

OpenStudy (stickman1998):

So length is = x?

OpenStudy (stickman1998):

And height = x-4?

OpenStudy (math&ing001):

Yep, correct !

OpenStudy (stickman1998):

Tysm!

OpenStudy (math&ing001):

Yvw!